OpenStudy (anonymous):

The product of two consecutive even integers is 40 less than 8 times their sum. Find the two integers.

OpenStudy (anonymous):

n(n+1) +40= 8(n+n+1) n2+n +40 =16n +8 n2-15n+32 =0 solve for n